# Machine learning for option pricing:∑an empirical investigation of network architectures

### v2

@@ Abstract.
We consider the supervised learning problem of learning the price of an option or the implied volatility given appropriate input data (model parameters) and corresponding output data (option prices or implied volatilities). The majority of articles in this literature considers a (plain) feed forward neural network architecture in order to connect the neurons used for learning the function mapping inputs to outputs. In this article, motivated by methods in image classification and recent advances in machine learning methods for PDEs, we investigate empirically whether and how the choice of network architecture affects the accuracy and training time of a machine learning algorithm. We find that the generalized highway network architecture achieves the best performance, when considering the mean squared error and the training time as criteria, within the considered parameter budgets for the Black–Scholes and Heston option pricing problems. Considering the transformed implied volatility problem, a simplified DGM variant achieves the lowest error among the tested architectures. We also carry out a capacity-normalised comparison for completeness, where all architectures are evaluated with an equal number of parameters. Finally, for the implied volatility problem, we additionally include experiments using real market data.
@@ 1. Introduction
The empirical results show that the more advanced network architectures consistently outperform the MLP architecture, and should thus be preferred for the computation of option prices and implied volatilities. More specifically, we find that for option pricing problems, where we focus on the Black–Scholes and the Heston model, the generalized highway network architecture outperforms all other variants, when considering the mean squared error and the computational time as criteria. Moreover, we find that for the computation of the implied volatility, after a necessary transformation, a variant of the DGM architecture outperforms all other variants, when considering again the mean squared error and the computational time as criteria. See also Van Mieghem [30].
The remainder of this paper is organized as follows: in Section 2 we briefly introduce the models used in this article, and in Section 3 we outline the basics of neural networks. In Section 4 we describe the different network architectures that will be used throughout, namely residual networks, highway and generalized highway networks, as well as the DGM network and its two variants. In Section 5 we revisit the main problems and describe the method used for generating sample data. In Section 6 we present the results of the empirical analysis performed, where the different network architectures are compared, with main criteria being accuracy and computational time. In Section 7 we perform a capacity-normalised comparison in which all architectures are evaluated with an equal number of parameters. Finally, Section 8 contains a synopsis and some concluding remarks from the empirical analysis.
@@ 2. Models
We will first consider the Black and Scholes [4] model for the evolution of the asset price $S$, i.e.
@@
see e.g. Gatheral [14, Ch. 1]. The inverse problem in (2.3) does not admit a closed form solution and needs to be solved numerically, using methods such as the Newton–Raphson or Brent; see e.g. Liu et al. [22] for more details. The Black–Scholes model assumes that the volatility is constant, however the implied volatility computed from real market option prices exhibits a so-called smile or skew shape; see e.g. Figure 2.1 for a visualization.
@@
A popular model that captures the behaviour of market data is the Heston stochastic volatility model, cf. Heston [19]. The dynamics of the asset price in this model are provided by the following SDEs
@@ 3. Fundamentals of neural networks
In this section we briefly recall the basic building blocks of feedforward neural networks, fix notation, and introduce the multilayer perceptron (MLP) architecture used as a reference throughout the paper.
@@
A layer is a collection of perceptrons that share the same input. Let a layer have $n$ units, and collect their outputs into a vector $\mathbf{y}\in\mathbb{R}^{n}$. With a weight matrix $W\in\mathbb{R}^{n\times d}$ and bias vector $\mathbf{b}\in\mathbb{R}^{n}$, we can write
@@
Given a target output $\mathbf{y}$ and a neural network $H(\mathbf{x};\theta)$ with parameters $\theta$ (all weights and biases), we measure the discrepancy by a loss function $J$, e.g. the mean squared error,
$J(\theta):=\|H(\mathbf{x};\theta)-\mathbf{y}\|.$ (3.3)
@@
with learning rate $\alpha>0$. In practice, the training data are partitioned into mini-batches; the update (3.4) is applied on each batch, and one full pass over all batches is referred to as an epoch. We initialise all networks using either the Glorot [16] or He [17] schemes.
@@ 4. Network variations
We now describe the network architectures used in our experiments. All of them are based on the fully-connected MLP structure from Section 3, and differ only in how each layer transforms its input and how information is propagated across layers. Throughout this section we follow the notation of Srivastava et al. [28].
@@ 4.2. Highway network
In the approximation of complex functions using neural networks, the depth of the network is proven to be a significant factor in its success, see e.g. Telgarsky [29]. However, training deeper networks presents additional bottlenecks, such as the vanishing gradients problem. The intuition behind highway networks is to allow for unimpeded information flow across the layers; see Srivastava et al. [28]. Indeed, from equation (4.3) we can see that for the boundary values of $T$ we have
@@
where $I$ denotes the identity matrix. Thus, the transform gate $T$ allows the highway layer to vary its behavior between a standard MLP layer and an identity mapping, leaving $\mathbf{x}$ unaffected.
@@ 4.3. Generalized highway network
One can further generalize highway networks by removing the convexity requirement in the outcome. Let us define an additional non-linear transformation alongside $T(\mathbf{x},W_{T})$, which we call $C(\mathbf{x},W_{C})$, such that the outcome takes the form
@@
We call $T$ and $C$ the transform and carry gate, respectively, as they regulate the amount of information passed into the next layer from the transformed and original input. This gating mechanism allows for information to flow along the layers of the network without degradation. Highway networks are obviously a special case of generalized highway networks, once we set C=1TC=1-T. A visual representation of this layer is given in Figure 4.6.
@@ 4.4. DGM network
The DGM network [27] was proposed in the context of the Deep Galerkin Method for high-dimensional PDEs; see also Al-Aradi et al. [2]. Its layer structure is related to generalized highway networks but incorporates additional nonlinear blocks and an explicit dependence on the original input $\mathbf{x}$ at each layer.
@@ 4.4.1. DGM layer
A DGM layer takes as input the feature vector $\mathbf{x}$ and the output $S^{\mathrm{old}}$ of the previous layer (or an initial affine transform of $\mathbf{x}$). From these, it computes four intermediate vectors $Z,G,R,H$ and produces the new state $S^{\mathrm{new}}$ via
@@
    The vector $H$ is obtained by composing multiple nonlinear transformations, which can be interpreted as a small “subnetwork” inside the layer.
@@
As a consequence, a DGM layer contains substantially more parameters than an MLP layer and incurs higher computational cost, but offers increased flexibility for approximating functions with sharp local features.
@@ 4.4.2. Overall DGM architecture
which sets the dimension of the subsequent DGM layers. An output layer then maps the final state to the desired prediction. The overall architecture is summarised in Figure 4.8.
@@ 4.6. No-Recurrence DGM network
This reduces the parameter count and computational cost per layer while preserving some of the additional flexibility of the original DGM construction.
@@ 5. Problems revisited and data sampling
In this section, we take another look at the problems considered in this work and present the form that will be used in the empirical analysis. We also discus the methods for generating data and the parameter sets used for the training of the neural networks.
@@ 5.3. Data sampling
We treat each problem discussed above as a supervised learning problem. In other words, we require inputs and outputs which we can feed to the network and calculate the loss function. We will follow the configuration outlined in Liu et al. [22] for all three problems. We will generate a total of $1$ million samples for each problem, which we split into an $80/20$ training and validation set. We will generate an additional $100$ thousand samples used for evaluation. When sampling from the space of input parameters, we can either define a joint distribution over the entire domain, or sample each variable separately. Both Liu et al. [22] and Papazoglou-Hennig [24] opt for Latin hypercube sampling (LHS), in which values are sampled from a joint distribution, resulting in a better representation of the parameter space; cf. McKay et al. [23]. Thus, in order to remain consistent with previous research, we use LHS in this work to generate the data as well.
Starting from the Black–Scholes pricing problem, we sample points using the parameters shown in Table 5.11, and use the analytical solution in equation (5.1) to obtain the output labels. We use the same sampling parameters as in [22, 24], allowing us to compare the network performances to results in both works at a later stage.
@@
The Heston pricing problem obviously requires additional parameters, whose range is listed in Table 5.12. The range for these parameters is selected again in accordance with Liu et al. [22].
@@ 6. Empirical results
In this section, we train neural networks using the architectures discussed in Sections 3 and 4, and apply them to the three different problems outlined in Sections 2 and 5. We then compare the empirical performance of these networks against each other, with the aim of identifying an optimal network architecture for the solution of these and related problems. The implementation for every network was done using Google’s TensorFlow library for Python [1] and the empirical analysis was carried out utilizing the Delft High Performance Computing Center [9] facilities.
@@ 6.1. Multilayer perceptrons
We start by training a set of MLPs ranging from a very small network to a large network, with similar configurations. The goal is to evaluate the performance of this network architecture, as well as to determine whether the size of the network influences the performance over a fixed range of epochs. To this end, we define a set of twelve networks with configurations listed in Table 6.15. We fix the learning rate according to the optimal value found from the study conducted in Liu et al. [22]. The batch size is chosen as large as possible while keeping calculations on a CPU fast. We found that a batch size of $64$ works well for the supervised learning problems discussed in this work. Furthermore, each network in this work is trained for $200$ epochs on the entire training dataset. We found out that, in most cases, increasing the amount of training epochs is beneficial. However, due to the large amount of network architectures that must be trained, we decided to terminate training after $200$ epochs.
@@ 6.1.2. Heston model
The results of the various MLPs trained and evaluated on the Heston pricing problem are presented in Figure 6.19. Therein we observe that the results for the Heston pricing problem using the MLP network architectures align qualitatively with those for the Black–Scholes problem. Since both problems are very similar in nature and attempt to approximate the same quantity, the similarity in results is rather expected.
@@
Indeed, the results from both the Black–Scholes as well as the Heston pricing problem exhibit, aside from the magnitude of the MSEs, many similarities; compare Figures 6.16 and 6.19. The MSEs and training times for the Heston model can be found in Table 6.20. The difference in MSEs is not surprising, as the Heston model has 8 parameters compared to the 4 parameters in the Black–Scholes model.
@@ 6.1.3. Implied volatility
The identical performance of the networks during training is visualized in Figure 6.22. The results we find align with the errors found in [24], in which the best performing MLP network containing 202,201202,\!201 parameters performed worse on the test data than the considerably smaller networks from Figure 6.21. Table 6.23 lists all the networks along with the training times, amount of parameters and MSE for the implied volatility problem. The results indicate that the networks are dealing with convergence issues, preventing them to optimize further, regardless of their size and architecture.
@@
Figure 6.24 visualizes the performances of the MLPs on the transformed implied volatility problem. We notice that on the transformed problem, the MSE is decreasing as a function of the network size, similarly to what we found for the option pricing problems before. The reduction in MSE is not as consistent as in the previous problems, indicating difficult convergence during training, possibly still due to a steep gradient not being entirely removed by the transformation. The results are considerably better than for the ‘default’ implied volatility problem, using only a simple transformation. Table 6.25 shows the results of all MLPs alongside their parameters and training times.
@@
To quantify the practical impact of the log-transformation we compare the performance of identical architectures on the raw and transformed implied volatility datasets. Using the 2-layer MLP with 150 nodes per layer, the raw formulation (Table 6.23) yields a test MSE of $7.1\times 10^{-4}$, whereas the transformed formulation (Table 6.25) reduces the error to $1.5\times 10^{-6}$. This improvement by almost three orders of magnitude is achieved without changing the architecture, hyperparameters, or training regime. The substantial gain in accuracy confirms empirically that the time–value and log–transform significantly improve numerical conditioning and optimisation stability, consistent with the gradient analysis in Section 5.2.
@@ 6.2.1. Black–Scholes model
Figure 6.27 visualizes the training times (blue) and MSE on the test set (red) for each highway network. Both the highway and generalized highway networks perform better than their MLP counterpart with an equal amount of nodes. More impressive, however, is the fact that both networks required barely more training time than the small MLP, but scored comparably to the large MLP on the test set. This means that the highway architecture improves convergence on the Black–Scholes problem significantly. The difference in convergence speed can also be seen during training, as shown in Figure 6.28. In this case, training losses are a good indicator for performance on the test set, likely because the performance difference for both networks is so large. The loss during training is roughly one order of magnitude smaller for the generalized highway model compared to the MLP, whereas the training time is only $23\%$ longer; cf. Table 6.29.
@@ 6.2.2. Heston model
Figure 6.30 visualizes the performance of the highway networks for the Heston pricing problem. We observe again a similar behavior as in the Black–Scholes pricing problem; compare with Figure 6.27. Indeed, the highway and generalized highway networks perform comparably to the large MLP in terms of MSE, while requiring the training time of the small MLP. A major difference is the performance of the residual network, which reduced the MSE by $33\%$ compared to the MLP on the Black–Scholes problem; cf. Table 6.29. On the Heston pricing problem, we see no increases in performance. A possible explanation for this could be that the Heston pricing problem might be too involved for the residual architecture to make an improvement.
@@
The highway and generalized highway networks do, in fact, improve the performance significantly compared to the MLP networks, with the generalized highway network performing almost as well as the largest MLP, while requiring only $31\%$ of the computational time. Table 6.31 lists the reductions that each highway variation contributes to the MSE. We notice that the improvement of the generalized highway network on the Heston problem is less significant than on the Black–Scholes problem, where the generalized highway network was the best performing network. In Section 6.3 we compare the highway networks to their DGM counterpart on the Heston pricing problem, and consider equally sized networks in terms of parameters.

### v3

@@ Abstract.
We consider the supervised learning problem of learning the price of an option or the implied volatility given appropriate input data (model parameters) and corresponding output data (option prices or implied volatilities). The majority of articles in this literature considers a (plain) feedforward neural network architecture in order to connect the neurons used for learning the function mapping inputs to outputs. In this article, motivated by methods in image classification and recent advances in machine learning methods for PDEs, we investigate empirically whether and how the choice of network architecture affects the accuracy and training time of a machine learning algorithm. We find that highway-type network architectures achieve the best performance, when considering the mean squared error and the training time as criteria, within the considered parameter budgets for the Black–Scholes and Heston option pricing problems. Considering the transformed implied volatility problem, a simplified DGM variant achieves the lowest error among the tested architectures. We also carry out a capacity-normalised comparison for completeness, where all architectures are evaluated with an equal number of parameters. Finally, for the implied volatility problem, we additionally include experiments using real market data.
@@ 1. Introduction
The empirical results show that the more advanced network architectures consistently outperform the MLP architecture, and should thus be preferred for the computation of option prices and implied volatilities. More specifically, we find that for option pricing problems, where we focus on the Black–Scholes and the Heston model, highway-type network architectures outperform all other variants, when considering the mean squared error and the computational time as criteria. Moreover, we find that for the computation of the implied volatility, after a necessary transformation, a variant of the DGM architecture outperforms all other variants, when considering again the mean squared error and the computational time as criteria. See also Van Mieghem [30].
The remainder of this paper is organized as follows: in Section 2 we briefly introduce the models used in this article, and in Section 3 we outline the basics of neural networks. In Section 4 we describe the different network architectures that will be used throughout, namely residual networks, highway and generalized highway networks, as well as the DGM network and its two variants. In Section 5 we revisit the main problems and describe the method used for generating sample data. In Section 6 we present the results of the empirical analysis performed, where the different network architectures are compared, with main criteria being accuracy and computational time. Moreover, we perform a capacity-normalised comparison in which all architectures are evaluated with an equal number of parameters, and also apply our results to real market data. Finally, Section 7 contains a synopsis and some concluding remarks from the empirical analysis.
@@ 2. Models
We first consider the Samuelson / Black and Scholes [4] model for the evolution of the asset price $S$, i.e.
@@
see e.g. Gatheral [14, Ch. 1]. The inverse problem in (2.3) does not admit a closed form solution and needs to be solved numerically, using methods such as the Newton–Raphson or Brent; see e.g. Liu et al. [22] for more details. The Black–Scholes model assumes that the volatility is constant; however the implied volatility computed from real market option prices exhibits a so-called smile or skew shape; see e.g. Figure 2.1 for a visualization.
@@
A popular model that captures the behavior of market data is the Heston stochastic volatility model, cf. Heston [19]. The dynamics of the asset price in this model are provided by the following SDEs
@@ 3. Fundamentals of neural networks
In this section, we briefly recall the basic building blocks of feedforward neural networks, fix notation, and introduce the multilayer perceptron (MLP) architecture used as a reference throughout the paper.
@@
A layer is a collection of perceptrons that share the same input. Let a layer have $n$ units, and collect their outputs into a vector $\mathbf{y}\in\mathbb{R}^{n}$. Using a weight matrix $W\in\mathbb{R}^{n\times d}$ and bias vector $\mathbf{b}\in\mathbb{R}^{n}$, we can write
@@
Consider a target output $\mathbf{y}$ and a neural network $H(\mathbf{x};\theta)$ with parameters $\theta$ (all weights and biases). We measure the discrepancy by a loss function $J$, e.g. the mean squared error,
$J(\theta):=\|H(\mathbf{x};\theta)-\mathbf{y}\|^{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2}.$ (3.3)
@@
with learning rate $\alpha>0$. In practice, the training data are partitioned into mini-batches; the update (3.4) is applied on each batch, and one full pass over all batches is referred to as an epoch. We initialize all networks using either the Glorot [16] or the He [17] schemes.
@@ 4. Network variations
Let us now describe the network architectures used in our experiments. All of them are based on the fully-connected MLP structure from Section 3, and differ only on how each layer transforms its input, and how information is propagated across layers. Throughout this section, we follow the notation of Srivastava et al. [28].
@@ 4.2. Highway network
In the approximation of complex functions using neural networks, the depth of the network is proven to be a significant factor in its success, see e.g. Telgarsky [29]. However, training deeper networks presents additional bottlenecks, such as the vanishing gradients problem. The intuition behind highway networks is to allow for unimpeded information flow across layers; see Srivastava et al. [28]. Indeed, from equation (4.3) we can see that for the boundary values of $T$ we have
@@
where $I$ denotes the identity matrix. Thus, the transform gate $T$ allows the highway layer to vary its behaviour between a standard MLP layer and an identity mapping, leaving $\mathbf{x}$ unaffected.
@@ 4.3. Generalized highway network
One can further generalize highway networks by removing the convexity requirement in the outcome. Let us define an additional non-linear transformation alongside $T(\mathbf{x},W_{T})$, which we denote by $C(\mathbf{x},W_{C})$, such that the outcome takes the form
@@
We call $T$ and $C$ the transform and carry gate, respectively, as they regulate the amount of information passed into the next layer from the transformed and original input. This gating mechanism allows information to flow along the layers of the network without degradation. Highway networks are obviously a special case of generalized highway networks, once we set C=1TC=1-T. A visual representation of this layer is given in Figure 4.4.
@@ 4.4. DGM network
The DGM network architecture was proposed in the context of the Deep Galerkin Method for the solution of PDEs; see Sirignano and Spiliopoulos [27] and also Al-Aradi et al. [2]. This layer structure is related to generalized highway networks but incorporates additional nonlinear blocks and an explicit dependence on the original input $\mathbf{x}$ at each layer.
@@ 4.4.1. DGM layer
A DGM layer takes as input the feature vector $\mathbf{x}$ and the output $S^{\mathrm{old}}$ of the previous layer (or an initial affine transform of $\mathbf{x}$). Using these, it computes four intermediate vectors $Z,G,R,H$ and produces the new state $S^{\mathrm{new}}$ via
@@
    The vector $H$ is obtained by composing multiple nonlinear transformations, which can be interpreted as a small “subnetwork” within the layer.
@@
These two aspects will be further explored in the next subsections, by introducing and testing new architectures that stress these points. As a consequence, a DGM layer contains substantially more parameters than an MLP or a highway layer and incurs higher computational cost, but offers increased flexibility for approximating functions with sharp local features.
@@ 4.4.2. Overall DGM architecture
which sets the dimension of the subsequent DGM layers. An output layer then maps the final state to the desired prediction. The overall architecture is summarized in Figure 4.6.
@@ 4.6. No-Recurrence DGM network
This reduces the parameter count and computational cost per layer, while preserving some of the additional flexibility of the original DGM construction.
@@ 5. Problems revisited and data sampling
In this section, we take another look at the problems considered in this work and present the form that will be used in the empirical analysis. We also discuss the methods for generating data and the parameter sets used for training the neural networks.
@@ 5.3. Data sampling
We treat each problem discussed above as a supervised learning problem. In other words, we require inputs and outputs which we can feed to the network and calculate the loss function. We will follow the configuration outlined in Liu et al. [22] for all three problems. We will generate a total of $1$ million samples for each problem, which we divide into an $80/20$ training and validation set. We will generate an additional $100$ thousand samples used for evaluation. When sampling from the space of input parameters, we can either define a joint distribution over the entire domain, or sample each variable separately. Liu et al. [22], as well as Papazoglou-Hennig [24] and Van Mieghem [30], opt for Latin hypercube sampling (LHS), in which values are sampled from a joint distribution, resulting in a better representation of the parameter space; cf. McKay et al. [23]. Thus, in order to remain consistent with previous research, we use LHS in this work to generate the data as well.
Starting from the Black–Scholes pricing problem, we sample points using the parameters shown in Table 5.1, and use the analytical solution in equation (5.1) to obtain the output labels. We use the same sampling parameters as in [22, 24], allowing us to compare the network performances with the results of both works at a later stage.
@@
The Heston pricing problem obviously requires additional parameters, whose range is listed in Table 5.2. The range for these parameters is selected again according to Liu et al. [22].
@@ 6. Empirical results
In this section, we train neural networks using the architectures discussed in Sections 3 and 4, and apply them to the three different problems outlined in Sections 2 and 5. We then compare the empirical performance of these networks against each other, with the aim of identifying an optimal network architecture for the solution of these and related problems. In addition, we test the method for the computation of implied volatility using real data, and also briefly compare this with existing methods. The implementation for every network was done using Google’s TensorFlow library for Python [1] and the empirical analysis was carried out using the Delft High Performance Computing Center [9] facilities.
@@ 6.1. Multilayer perceptrons
We start by training a set of MLPs ranging from a very small network to a large network, with similar configurations. The goal is to evaluate the performance of this network architecture, as well as to determine whether the size of the network influences the performance over a fixed range of epochs. To this end, we define a set of twelve networks with configurations listed in Table 6.1. We fix the learning rate according to the optimal value found from the study conducted in Liu et al. [22]. The batch size is chosen as large as possible while keeping the calculations on a CPU fast. We found that a batch size of $64$ works well for the supervised learning problems discussed in this work. Furthermore, each network in this work is trained for $200$ epochs on the entire training dataset. We found that, in most cases, increasing the number of training epochs is beneficial. However, due to the large amount of network architectures that must be trained, we decided to terminate training after $200$ epochs.
@@ 6.1.2. Heston model
The results of the various MLPs trained and evaluated on the Heston pricing problem are presented in Figure 6.5. Therein we observe that the results for the Heston pricing problem using the MLP network architectures align qualitatively with those for the Black–Scholes problem. Since both problems are very similar in nature and attempt to approximate the same quantity, the similarity in the results is rather expected.
@@
Indeed, the results from both the Black–Scholes as well as the Heston pricing problem exhibit, aside from the magnitude of the MSEs, many similarities; compare Figures 6.2 and 6.5. The MSEs and training times for the Heston model can be found in Table 6.6. The difference in MSEs is not surprising, as the Heston model has 8 parameters compared to the 4 parameters in the Black–Scholes model.
@@ 6.1.3. Implied volatility
The identical performance of the networks during training is visualized in Figure 6.8. The results we find align with the errors found in [24], in which the best performing MLP network containing 202,201202,\!201 parameters performed worse on the test data than the considerably smaller networks from Figure 6.7. Table 6.9 lists all the networks along with the training times, number of parameters and MSE for the implied volatility problem. The results indicate that the networks are dealing with convergence issues that prevent them from further optimization, regardless of their size and architecture.
@@
Figure 6.10 visualizes the performance of the MLPs on the transformed implied volatility problem. We notice that on the transformed problem, the MSE is decreasing as a function of the network size, similarly to what we found for the option pricing problems before. The reduction in MSE is not as consistent as in the previous problems, indicating difficult convergence during training, possibly still due to a steep gradient not being entirely removed by the transformation. The results are considerably better than for the ‘default’ implied volatility problem, using only a simple transformation. Table 6.11 shows the results of all MLPs alongside their parameters and training times.
@@
In order to quantify the practical impact of the log-transformation we compare the performance of identical architectures on the default and transformed implied volatility datasets. Using the 2-layer MLP with 150 nodes per layer, the default formulation (Table 6.9) yields a test MSE of $7.1\times 10^{-4}$, whereas the transformed formulation (Table 6.11) reduces the error to $1.5\times 10^{-6}$. This improvement by almost three orders of magnitude is achieved without changing the architecture, hyperparameters, or training regime. The substantial gain in accuracy confirms empirically that the time-value and log-transform significantly improve numerical conditioning and optimization stability, consistent with the gradient analysis in Section 5.2.
@@ 6.2.1. Black–Scholes model
Figure 6.13 visualizes the training times (blue) and MSE on the test set (red) for each highway network. Both the highway and generalized highway networks perform better than their MLP counterpart with an equal number of nodes. More impressive, however, is the fact that both networks required barely more training time than the small MLP, but scored comparably to the large MLP on the test set. This means that the highway architecture improves convergence on the Black–Scholes problem significantly. The difference in convergence speed can also be seen during training, as shown in Figure 6.14. In this case, training losses are a good indicator for performance on the test set, likely because the performance difference for both networks is so large. The loss during training is roughly one order of magnitude smaller for the generalized highway model compared to the MLP, whereas the training time is only $23\%$ longer; cf. Table 6.15.
@@ 6.2.2. Heston model
Figure 6.16 visualizes the performance of the highway networks for the Heston pricing problem. We observe again a similar behaviour as in the Black–Scholes pricing problem; compare with Figure 6.13. Indeed, the highway and generalized highway networks perform comparably to the large MLP in terms of MSE, while requiring the training time of the small MLP. A major difference is the performance of the residual network, which reduced the MSE by $33\%$ compared to the MLP on the Black–Scholes problem; cf. Table 6.15. On the Heston pricing problem, we do not observe an increase in performance. A possible explanation for this could be that the Heston pricing problem might be too involved for the residual architecture to make an improvement.
@@
The highway and generalized highway networks do, in fact, significantly improve performance compared to the MLP networks, with the generalized highway network performing almost as well as the largest MLP, while requiring only $31\%$ of the computational time. Table 6.17 lists the reductions that each highway variation contributes to the MSE. We notice that the improvement of the generalized highway network on the Heston problem is less significant than on the Black–Scholes problem, where the generalized highway network was the best performing network. In Section 6.3 we compare the highway networks to their DGM counterpart on the Heston pricing problem, and consider equally sized networks in terms of parameters.